OpenStudy (anonymous):

How do you calculate the number of mole: 3.011x10^23 atoms of plutonium

OpenStudy (btaylor):

1 mole of a substance is 6.022x10^23 atoms/molecules. \[3.011 \times 10^{23} atoms \times \frac{1 mol}{6.022 \times 10^{23} atoms} = \frac{1}{2} mol\]

OpenStudy (btaylor):

You calculate it by multiplying the number of atoms you have by the ratio moles to atoms.